Neural Computation, Volume 1, Number 1

CONTENTS

Reviews Review of Neural Networks for Speech Recognition

Richard P. Lippmann 1

Letters Modular Construction of Time-Delay Neural Networks for Speech Recognition

Alex Waibel 39

A Silicon Model of Auditory Localization John Lazzaro and Carver A. Mead 47

Criteria for Robust Stability In A Class Of Lateral Inhibition Networks Coupled Through Resistive Grids

John L. Wyatt, Jr., and David L. Standley 58

Two Stages of Curve Detection Suggest Two Styles of Visual Computation Steven W. Zucker, Allan Dobbins, and Lee Iverson 68

Part Segmentation for Object Recognition Alex Pentland 82

Computing Optical Flow in the Primate Visual System H. Taichi Wang, Bimal Mathur, and Christof Koch 92

A Multiple-Map Model for Pattern Classification Alan Rojer and Eric Schwartz 104

A Control Systems Model of Smooth Pursuit Eye Movements with Realistic Emergent Properties

R. J. Krauzlis and S. G. Lisberger 116

The Brain Binds Entities and Events by Multiregional Activation from Convergence Zones

Antonio R. Damasio 123

Product Units: A Computationally Powerful and Biologically Plausible Extension to Backpropagation Networks

Richard Durbin and David E. Rumelhart 133

Deterministic Boltzmann Learning Performs Steepest Descent in Weight-Space

Geoffrey E. Hinton 143

What Size Net Gives Valid Generalization?

Eric B. Baum and David Haussler 151